OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (anonymous):

\[ (x-h)^2 + (y-k)^2 = r^2 \]

OpenStudy (anonymous):

\[(x-h)^2 + (y-k)^2 = r^2\\ (x-2)^2 + (y-(-1))^2 = 8^2\\ (x-2)^2 + (y+1))^2 = 64 \]

OpenStudy (anonymous):

Thank you!